# Forms of Understanding for XAI-Explanations

### v2

@@ Abstract
[ubi]organization=Bielefeld University, addressline=Universitätsstraße 25, city=Bielefeld, postcode=33602, country=Germany
@@
[upb]organization=Paderborn University, addressline=Warburger Straße 100, city=Paderborn, postcode=33098, country=Germany
@@
[trr]organization=SFB/Transregio 318 Constructing Explainability, city=Paderborn & Bielefeld, country=Germany
@@ 1 Introduction
Due to recent technological developments, explainability has become a widely discussed topic in computer science and artificial intelligence, giving rise to a new subfield called Explainable Artificial Intelligence (or XAI for short). It is also a topic of interest in other disciplines, such as cognitive science, psychology, (psycho-)linguistics, philosophy and education. The focus of this article is to provide a theoretical concept of understanding as a result of explanation. From this perspective, understanding is the goal of a successful explanation [53, p. 229] and “recipients of explanations, if the explanations are at all successful, are expanding their understanding in real time”. Understanding has implicit knowledge parts [96, p. 97] regarding its procedural and non-reflective components. As explanations progress, the level of understanding increases, but understanding always remains incomplete [67] and is not always factual [70], among other reasons because explanations are often shallow, depending on the knowledge and motivation of the explainer and the explainee [96], and they have “relatively unclear end states” [53, p. 242]. Thus, while there is thus a broad agreement that the goal of explanation is understanding, we know little about how understanding can be conceptualized, what forms of understanding can occur, and, finally, how it can be measured. The aim of this article is to bring together insights from several disciplines and to develop a concept based on their findings. We will propose a four-field model of understanding that provides conceptual clarification of related terms and concepts, and thus distinguishes between different forms and levels of understanding. Our goal is to develop a model that can be used as a heuristic to capture, analyze, and even measure different forms of understanding in the context of XAI and beyond, and that can be operationalized in interactive explanation systems.
@@
The starting point of our considerations is a view of an XAI-explanation as a co-constructive, interactive process to which both humans and XAI or interactive explanation systems contribute [82]. The goal of an explanation – i.e., the form and/or level of understanding it aims to achieve – depends on the goal of the explainer as well as contextual and situational contingencies [82]. Rather than focusing solely on the narrow case of explainability, i.e., understanding the inner workings of an AI system using XAI methods, we consider XAI to have a broader scope of explaining more generally (e.g., explaining a programming language concept to a novice computer programmer, a grammar rule to a second language learner, or a board game to someone learning to play it) fostering understanding and agency of humans. Accordingly, we adopt a broad notion of ‘understanding’ as implemented in and underlying everyday interactions and experiences with algorithmic technology in socio-technical AI systems. In contrast to other research on explanation, we do not focus on scientific explanations but on ‘everyday explanations’ because they show a greater variety with regard to the knowledge they mediate [53]. While scientific explanations are regularly characterized by why-questions [e.g., 22], and thus aim primarily to convey causal knowledge, everyday explanations often follow other kinds of questions, such as what, where, how does it work, etc. [56]. Moreover, the quality of everyday explanations is more difficult to define – they can be very simple, incomplete, inaccurate, or, in contrast, quite sophisticated [50, 69].
@@
The above-mentioned disciplines have dealt with the problem of understanding to varying degrees. While understanding is a widely discussed topic in philosophy, it is not a topic in experimental psychology nor is there a scientific concept of it. Likewise, ‘understanding’ is hardly considered a research topic in its own right in many disciplines [but see, e.g, 99]. Nevertheless, this article will provide a synthesis of some interdisciplinary perspectives on ‘understanding’ as a result of everyday explanations, i.e., non-scientific explanations that occur in everyday contexts and can involve either humans or a human and a machine. The objective is to present a model of the different forms of understanding that can arise from them. This is highly relevant for tutoring contexts [e.g., 23, 67], classroom interaction [e.g., 57, 76, 65], as well as for XAI [82]. We would like to argue that a model that attempts to define and unify basic concepts is a useful tool across disciplines because it can provide a starting point for both further empirical analysis and theory development. While the proposed model is primarily concerned with cognitive aspects, we will also argue that embodied interaction should be included. With such a claim, we intend to expand the notion of understanding from a purely cognitive conception to one that includes bodily aspects. The model will be enriched by offering a theoretical account of the assumed dynamics of understanding by sketching how movement between the four fields can be conceptualized.
@@ 1.1 Basic concepts and first examples
In this paper, explaining is conceptualized as a social interactive process [66] in which an explainer (‘EX’) and an explainee (‘EE’) iteratively co-construct an explanans of an entity that is explained (the explanandum). This concept is based on everyday explanations which, in contrast to scientific explanations, are always interactive and usually embedded in a situational context. The explainer is an agent that can either be a human or, as it is the case in XAI-explanations, an interactive explanation system. Both partners are actively involved in the explanatory process through mutual monitoring and scaffolding [82] and proposals have been made how to model such activities for interactive explanation systems [80, 15, 18]. The goal of an explanatory process is for the explainees to understand the explanandum with regard to their situational objectives or purposes; the form and depth of understanding, however, can vary according to the explainee’s goals as well as situational and contextual contingencies. This concept of explaining that emphasizes the process rather than the (normatively fixed) outcome differs from the view often held in science that explanations are products with formalized normative forms, and comply to rules of logic [45]. Everyday explanations also include causal, procedural or conceptual relations [78], but usually do not follow forms or rules of natural science; neither do they necessarily aim at formulating universal (scientific) laws. Instead, everyday explanations are more tailored to the explainee’s situational objectives and needs. They are therefore more context-dependent than scientific explanations and correspondingly more flexible in terms of form and content [cf., 66].
@@
In this article, we will focus on ‘understanding’ as knowledge resulting from a given explanation. This should not mean that explanations always lead to understanding; indeed, they might lead to a false impression of understanding, or they may lead to no understanding at all, regardless of the explanation’s quality [48]. In fact, the connection between explanations and understanding is still of debate in philosophy. While some argue that understanding is conceptually prior to explanations, others claim that understanding is in fact knowing what a good explanation is, and the differences between understanding and explanatory knowledge are only minor [54]. While this is an interesting debate in itself, our focus here is different, namely to present a heuristics for forms of understanding which can arise as a result of successful knowledge transfer through everyday as well as XAI-explanations. We will thus not go into the relation between the two any further; instead, we assume for theoretical convenience that explanations can lead to knowledge transfer and thus to understanding on the part of the explainee.
@@
Before presenting a conceptual model of understanding, we assume that the notion of knowledge should be clarified. As mentioned above, we follow the view that it is an explanations’ objective to construct and transmit knowledge [66, 78], and we assume that understanding something on the basis of an explanation means knowing something. It should be noted that there is a debate in philosophy about differences between ‘understanding’ and ‘knowledge’. Knowledge, it is argued, can be rather easy to acquire, and objects of knowledge can be isolated [45]. Understanding, in contrast, seems to be more difficult to get because it involves connecting isolated knowledge objects with each other and combining them into a structured whole [99]. For the purpose of this article, we neglect the debate about possible differences between understanding and knowledge and take the view that the differences are marginal at best. However, building on previous scholarly work on knowledge and learning, a social constructivist view underlines the importance of the subjective parts of knowledge that arise from different goals, prior knowledge, etc. [cf., 71]. If two people read the same book, they will not have the same knowledge afterwards. Consider, for example, the geocentric worldview with the earth at the center of the universe, which was socially shared ‘knowledge’. Knowledge therefore also has subjective components. This is especially the case when we talk about the knowledge that results from understanding everyday explanations.
@@
The different forms of understanding that an explainer is aiming for – or that are sought after by the explainee – are reflected in the heterogeneity that can be found especially in everyday explanations. In order to deal with this diversity, we have developed a model that considers two types of understanding as the result or goal of an explanation: comprehension and enabledness. In addition, the model conceives of these types as further differentiated by the level of understanding that can be achieved, which we label shallow and deep respectively [cf. 42, p. 127]). It should be noted that our model, like any model, is simplified [9] in order to provide a systematic examination of different forms.
@@
In this article, we will use two examples to illustrate understanding (see examples 1 and 2): In the following, we would like to explain the terms introduced above using example 1. It looks at understanding in a technical context, namely the use of a stopping criterion in an iteration loop of a computer program. More specifically, example 1 is an illustration to learn how to use while-loops in computer programming. While-loops are constructs of programming languages that execute the code within the construct as long as the while-condition attached to the construct holds. In the example, a variable counter is decremented by $1$ as long as its current value is greater than $1$. The specific application of the example is to print the well known ‘99 bottles’ song [^1] https://en.wikipedia.org/wiki/99_Bottles_of_Beer [19]. While-loops are known to cause problems for computer programming students [77] and could be a topic for an interactive explanation system in the domain of computer science education.
{mdframed}
[…]
5:$counter\leftarrow counter-1$
[…]
@@
of the program would entail that it will print out the ‘99 bottles’-song on the computer screen. Specifically, regarding the while-construct, shallow comprehension could involve understanding that the indented lines (3–7) are executed every time the while-loop is entered and that counter $>$ 0 in line 2 is the condition under which the while-loop is executed.
@@
{mdframed}
This example looks at a specific construction of natural language use, specifically the English present progressive tense. The present progressive is constructed using “the present tense form of be + lexical verb in -ing form” [20, p. 598] and is used for, among other things, “events in progress at the time of speaking”, “repeated events in temporary contexts”, “processes of change”, and “with adverbs of indefinite frequency” [20, p. 601f], i.e., it expresses aspect grammatically. These usage contexts may be intuitive to native English speakers, who may also be able to see the nuanced differences in usage compared to the present simple tense. However, learners of English as a second language, especially speakers of languages where there is no grammatical aspect (e.g., standard German), need to understand how and when to use the present progressive and may not even have the concepts (e.g., that progressivity is grammatically marked) to begin with.
@@
Section 3, will look at both examples in more detail.
@@ 2 Aspects of Understanding: Forms of knowledge
For our model of understanding, we follow primarily two conceptualizations: the types and qualities of knowledge as described in detail by de Jong and Ferguson-Hessler [51], and the taxonomies of learning from the field of education [‘Bloom’s taxonomy’; 6, 59].
de Jong and Ferguson-Hessler [51] were interested in knowledge in instruction and learning and especially as a prerequisite for problem solving. They distinguished and elaborated characteristics of different types of knowledge. In terms of understanding as a goal and the achievement of explanations, two types of knowledge are important: conceptual knowledge is “knowledge of facts, concepts, and principles” (p. 107), whereas procedural knowledge refers to “actions and manipulations” that transform one state into another. Thus, in our terminology, conceptual knowledge describes ‘knowing-that’, and procedural knowledge refers to ‘knowing-how’. In the context of XAI, we focus on cognitive forms of procedural knowledge, such as written division and programming. Actions and skills with a higher motor component, such as cycling or knitting, may require other forms of more embodied explanations (showing).
@@
Building on the taxonomy of qualities of knowledge also developed by de Jong and Ferguson-Hessler [51], we consider the following characteristics to be relevant for describing understanding:
@@
Both conceptual and procedural knowledge are described as more or less shallow or deep. Regarding conceptual knowledge, de Jong and Ferguson-Hessler [51, p. 111] illustrate the distinction with symbols or formulas on a shallow level vs. concepts and their relations on a deep level. Thus, deep comprehension most likely reflects highly interconnected and elaborated concepts. Shallow comprehension would correspond to little interconnectedness, but also to few concepts. In terms of procedural knowledge, there are rules or recipes at a shallow level in contrast to meaningful actions at a deep level.
Explainees may require different levels of understanding, and explanations should be tailored to those needs [96]. Thus, it is not always necessary – or even annoying – to give or receive a long and complex explanation that leads to deep understanding. Often the explainer is satisfied with a single action or a simple fact.
@@
Automation of knowledge is particularly important for understanding and therefore for explanation. Consider how difficult it is to explain simple skills like riding a bicycle. This is because there is a strong distinction between declarative conceptual knowledge and non-declarative compiled procedural knowledge. Compilation enables automatic processes. The term “knowledge compilation”, inspired by computer science, is used by Anderson [3] to describe how multiple cognitive steps are combined into a single routine. Compilation involves two stages: composition, where steps are combined, and proceduralization, where declarative knowledge becomes automatic, resulting in more fluid performance.
We conceive comprehension as involving conceptual knowledge, i.e., remembering facts about and experiences with various concepts such as theories and/or related routine actions. In contrast, enabledness is conceived as the ability to perform routine and/or complex actions (mentally and physically) in a competent and flexible manner. Furthermore, whereas comprehension is always ‘explicit knowledge’ that can be acquired through books or explicit verbal instruction, enabledness consists at least in part of ‘tacit knowledge’ that is not easily shared [62].
However, conceptual knowledge can also be compiled, e.g., rote learning of definitions or mnemonic sentences instead of deep comprehension. On the other hand, deep enabledness requires declarative, mostly deep comprehension. In addition to implicit learning of procedures as a basis for skills, e.g., native language grammar, there is explicit learning, e.g., adult learning of a foreign language. Explaining how to drive a car is a good example of how procedures are first explained in a declarative way, and then the steps are proceduralized and composed over the years [3]. de Jong and Ferguson-Hessler [51] explain automation by mentioning tacit knowledge that cannot be expressed. In order to give an explanation, the explainer needs non-tacit knowledge, also of procedures. In the following sections, this intertwining of understanding quality and type will be discussed in detail.
@@
Regarding modality, we added embodiment to the verbal and pictorial aspects given by de Jong and Ferguson-Hessler [51]. Embodied knowledge is conceptually similar to procedural knowledge [88, p. 150], and it can be better performed than explained verbally. Embodiment plays an important role in learning situations where new members of a (scientific) community are taught to become skilled practitioners. Visibly embodied displays (e.g., how to hold a pencil, how to hold a Munsell color chart next to a piece of dirt to describe its color) are part of the knowledge conveyed in such situations [41].
@@
In addition to this characterization of different types of knowledge and their manifold qualities, as a further conceptualization, taxonomies of learning objectives help to describe the types and levels of understanding [e.g., 59, 6]. Anderson and Krathwohl 6 (summarized in Anderson 5) systematize educational objectives in two dimensions, the structure of the knowledge and the cognitive processes that are to be done with the knowledge.
@@
The structure of knowledge aims at a differentiation of the above mentioned ‘types of knowledge’. It includes factual, conceptual, procedural, and metacognitive knowledge. While factual knowledge contains the basic elements of a domain (e.g., definitions), conceptual knowledge contains the structure of the basic elements, e.g., principles, models, and theories. Comparable to de Jong and Ferguson-Hessler [51], procedural knowledge aims at ‘how to do something’, such as skills and algorithms. The fourth form of knowledge, metacognition, will be neglected in our paper, because it is only of interest in special explanations regarding (own) cognition. Thus, the categories of knowledge are similar to the types of knowledge discussed by de Jong and Ferguson-Hessler [51] with the additional hint that there are very basic elements of knowledge. In terms of understanding, factual and conceptual knowledge aim at comprehension, while procedural knowledge, of course, aims at enabledness.
@@
The second dimension of the taxonomy of learning aims at increasing the levels of cognitive processes that can be done with knowledge, namely remembering, understanding (note that this is a different use of the term), applying, analyzing, evaluating, and creating [24]. This strengthens the aspect of the level of knowledge introduced by de Jong and Ferguson-Hessler [51]. There is a gain in complexity from low level cognitive processes like remembering to applying to high level processes like creating something new – resulting in a hierarchy from shallow to deep understanding. However, assessing the level of understanding makes it necessary to take a closer look at these processes. For example, whether creating something new is a high-level process could depend on what is meant by ‘new’ and the complexity of what is being created, e.g., a new sentence or a scientific article.
@@
Shallow Enabledness Deep Enabledness
Shallow Comprehension Sect. 3.1 Sect. 3.2
Deep Comprehension Sect. 3.3 Sect. 3.4
Figure 1: Types and levels of understanding as a four field diagram
@@
In summary, there are two forms of understanding – comprehension and enabledness. Both forms of understanding can be on different levels, from very shallow to very deep understanding. The four resulting combinations (cf. Figure 1) are intermediate steps or final achievements of explanations with the possible goal of agency, the ability to control and adapt the (digital) environment (Passey et al. 72; Section 3.4). The combinations will be characterized in the next section. On this basis, (in Section 4) we explain dynamics of comprehension in explanations, taking into account two relevant aspects: We consider differences between novices and experts in explanations and discuss how in the course of an explanation comprehension and enabledness are intertwined.

### v3

@@ Abstract
[ubi]organization=Bielefeld University, city=Bielefeld, country=Germany
@@
[upb]organization=Paderborn University, city=Paderborn, country=Germany
@@
[uwk]organization=University for Continuing Education Krems, city=Krems, country=Austria
[…]
[trr]organization=SFB/Transregio 318 ‘Constructing Explainability’, city=Paderborn & Bielefeld, country=Germany
@@ 1 Introduction
Due to recent technological developments, explainability has become a widely discussed topic in computer science and artificial intelligence, giving rise to a new subfield called Explainable Artificial Intelligence (or XAI for short). It is also a topic of interest in other disciplines, such as cognitive science, psychology, (psycho-)linguistics, philosophy and education. The focus of this article is to provide a theoretical concept of understanding as a result of explanation. From this perspective, understanding is the goal of a successful explanation (Keil, 2006, p. 229) and “recipients of explanations, if the explanations are at all successful, are expanding their understanding in real time”. Understanding has implicit knowledge parts (Wilson and Keil, 2000, p. 97) regarding its procedural and non-reflective components. As explanations progress, the level of understanding increases, but understanding always remains incomplete (Miyake, 1986) and is not always factual (Páez, 2019), among other reasons because explanations are often shallow, depending on the knowledge and motivation of the explainer and the explainee (Wilson and Keil, 2000), and they have “relatively unclear end states” (Keil, 2006, p. 242). Thus, while there is thus a broad agreement that the goal of explanation is understanding, we know little about how understanding can be conceptualized, what forms of understanding can occur, and, finally, how it can be measured. The aim of this article is to bring together insights from several disciplines and to develop a concept based on their findings. We will propose a four-field model of understanding that provides conceptual clarification of related terms and concepts, and thus distinguishes between different forms and levels of understanding. Our goal is to develop a model that can be used as a heuristic to capture, analyze, and even measure different forms of understanding in the context of XAI and beyond, and that can be operationalized in interactive explanation systems.
@@
The starting point of our considerations is a view of an XAI-explanation as a co-constructive, interactive process to which both humans and XAI or interactive explanation systems contribute (Rohlfing et al., 2021, 2025). The goal of an explanation – i.e., the form and/or level of understanding it aims to achieve – depends on the goal of the explainer as well as contextual and situational contingencies (Rohlfing et al., 2021). Rather than focusing solely on the narrow case of explainability, i.e., understanding the inner workings of an AI system using XAI methods, we consider XAI to have a broader scope of explaining more generally (e.g., explaining a programming language concept to a novice computer programmer, a grammar rule to a second language learner, or a board game to someone learning to play it) fostering understanding and agency of humans. Accordingly, we adopt a broad notion of ‘understanding’ as implemented in and underlying everyday interactions and experiences with algorithmic technology in socio-technical AI systems. In contrast to other research on explanation, we do not focus on scientific explanations but on ‘everyday explanations’ because they show a greater variety with regard to the knowledge they mediate (Keil, 2006). While scientific explanations are regularly characterized by why-questions (e.g., Chi et al., 1994), and thus aim primarily to convey causal knowledge, everyday explanations often follow other kinds of questions, such as what, where, how does it work, etc. (Klein, 2009). Moreover, the quality of everyday explanations is more difficult to define – they can be very simple, incomplete, inaccurate, or, in contrast, quite sophisticated (Johnson-Laird, 1983; Norman, 1988).
@@
The above-mentioned disciplines have dealt with the problem of understanding to varying degrees. While understanding is a widely discussed topic in philosophy, it is not a topic in experimental psychology nor is there a scientific concept of it. Likewise, ‘understanding’ is hardly considered a research topic in its own right in many disciplines (but see, e.g, Zagzebski, 2019). Nevertheless, this article will provide a synthesis of some interdisciplinary perspectives on ‘understanding’ as a result of everyday explanations, i.e., non-scientific explanations that occur in everyday contexts and can involve either humans or a human and a machine. The objective is to present a model of different forms of understanding that can arise from them. This is highly relevant for tutoring contexts (e.g., Chi et al., 2004; Miyake, 1986), classroom interaction (e.g., Koole, 2010; van de Pol et al., 2018; Martins and Gresse von Wangenheim, 2022), as well as for XAI (Rohlfing et al., 2021). We would like to argue that a model that attempts to define and unify basic concepts is a useful tool across disciplines because it can provide a starting point for both further empirical analysis and theory development. While the proposed model is primarily concerned with cognitive aspects, we will also argue that embodied interaction should be included. With such a claim, we intend to expand the notion of understanding from a purely cognitive conception to one that includes bodily aspects. The model will be enriched by offering a theoretical account of the assumed dynamics of understanding by sketching how movement between the four fields can be conceptualized.
@@ 1.1 Basic concepts and first examples
In this paper, explaining is conceptualized as a social interactive process (Miller, 2019, 2023) in which an explainer and an explainee iteratively co-construct an explanans of an entity that is explained (the explanandum). This concept is based on everyday explanations which, in contrast to scientific explanations, are always interactive and usually embedded in a situational context. The explainer is an agent that can either be a human or, as it is the case in XAI-explanations, an interactive explanation system. Both partners are actively involved in the explanatory process through mutual monitoring and scaffolding (Rohlfing et al., 2021) and proposals have been made how to model such activities for interactive explanation systems (Robrecht and Kopp, 2023; Booshehri et al., 2024; Buschmeier and Kopp, 2018). The goal of an explanatory process is for the explainees to understand the explanandum with regard to their situational objectives or purposes; the form and depth of understanding, however, can vary according to the explainee’s goals as well as situational and contextual contingencies. This concept of explaining that emphasizes the process rather than the (normatively fixed) outcome differs from the view often held in science that explanations are products with formalized normative forms, and comply to rules of logic (Grimm, 2021). Everyday explanations also include causal, procedural or conceptual relations (Quasthoff et al., 2017), but usually do not follow forms or rules of natural science; neither do they necessarily aim at formulating universal (scientific) laws. Instead, everyday explanations are more tailored to the explainee’s situational objectives and needs. They are therefore more context-dependent than scientific explanations and correspondingly more flexible in terms of form and content (cf., Miller, 2019).
@@
In this article, we will focus on ‘understanding’ as knowledge resulting from a given explanation. This should not mean that explanations always lead to understanding; indeed, they might lead to a false impression of understanding, or they may lead to no understanding at all, regardless of the explanation’s quality (Hempel, 1965). In fact, the connection between explanations and understanding is still of debate in philosophy. While some argue that understanding is conceptually prior to explanations, others claim that understanding is in fact knowing what a good explanation is, and the differences between understanding and explanatory knowledge are only minor (Khalifa, 2017). While this is an interesting debate in itself, our focus here is different, namely to present a heuristics for forms of understanding which can arise as a result of successful knowledge transfer through everyday as well as XAI-explanations. We will thus not go into the relation between the two any further; instead, we assume for theoretical convenience that explanations can lead to knowledge transfer and thus to understanding on the part of the explainee.
@@
Before presenting a conceptual model of understanding, we assume that the notion of knowledge should be clarified. As mentioned above, we follow the view that it is an explanations’ objective to construct and transmit knowledge (Miller, 2019; Quasthoff et al., 2017), and we assume that understanding something on the basis of an explanation means knowing something. It should be noted that there is a debate in philosophy about differences between ‘understanding’ and ‘knowledge’. Knowledge, it is argued, can be rather easy to acquire, and objects of knowledge can be isolated (Grimm, 2021). Understanding, in contrast, seems to be more difficult to get because it involves connecting isolated knowledge objects with each other and combining them into a structured whole (Zagzebski, 2019). For the purpose of this article, we neglect the debate about possible differences between understanding and knowledge and take the view that the differences are marginal at best. However, building on previous scholarly work on knowledge and learning, a social constructivist view underlines the importance of the subjective parts of knowledge that arise from different goals, prior knowledge, etc. (cf., Palincsar, 1998). If two people read the same book, they will not have the same knowledge afterwards. Consider, for example, the geocentric worldview with the earth at the center of the universe, which was socially shared ‘knowledge’. Knowledge therefore also has subjective components. This is especially the case when we talk about the knowledge that results from understanding everyday explanations.
@@
The different forms of understanding that an explainer is aiming for – or that are sought after by the explainee – are reflected in the heterogeneity that can be found especially in everyday explanations. In order to deal with this diversity, we have developed a model that considers two types of understanding as the result or goal of an explanation: comprehension and enabledness. In addition, the model conceives of these types as further differentiated by the level of understanding that can be achieved, which we label shallow and deep respectively (cf. Graesser and Person, 1994, p. 127). It should be noted that our model, like any model, is simplified (Bailer-Jones, 2009) in order to provide a systematic examination of different forms.
@@
In this article, we will use two examples to illustrate understanding (see examples 1 and 2): In the following, we would like to explain the terms introduced above using example 1. It looks at understanding in a technical context, namely the use of a stopping criterion in an iteration loop of a computer program. More specifically, example 1 is an illustration to learn how to use while-loops in computer programming. While-loops are constructs of programming languages that execute the code within the construct as long as the while-condition attached to the construct holds. In the example, a variable counter is decremented by $1$ as long as its current value is greater than $1$. The specific application of the example is to print the well known ‘99 bottles’ song [^1] https://en.wikipedia.org/wiki/99_Bottles_of_Beer (Byrd, 2010). While-loops are known to cause problems for computer programming students (Qian and Lehman, 2017) and could be a topic for an interactive explanation system in the domain of computer science education.
1: $counter\leftarrow 99$ 2: while $counter>0$ do 3:   output countercounter “ bottles of liquid on the wall,” 4:   output countercounter “ bottles of liquid;” 5:   output “If one of those bottles should happen to fall,” 6:   $counter\leftarrow counter-1$ 7:   output countercounter “ bottles of liquid on the wall.”
@@
of the program would entail that it will print out the ‘99 bottles’-song on the computer screen. Specifically, regarding the while-construct, shallow comprehension could involve understanding that the indented lines (3–7) are executed every time the while-loop is entered and that counter >> 0 in line 2 is the condition under which the while-loop is executed.
@@
This example looks at a specific construction of natural language use, specifically the English present progressive tense. The present progressive is constructed using “the present tense form of be + lexical verb in -ing form” (Carter and McCarthy, 2006, p. 598) and is used for, among other things, “events in progress at the time of speaking”, “repeated events in temporary contexts”, “processes of change”, and “with adverbs of indefinite frequency” (Carter and McCarthy, 2006, p. 601f), i.e., it expresses aspect grammatically. These usage contexts may be intuitive to native English speakers, who may also be able to see the nuanced differences in usage compared to the present simple tense. However, learners of English as a second language, especially speakers of languages where there is no grammatical aspect (e.g., standard German), need to understand how and when to use the present progressive and may not even have the concepts (e.g., that progressivity is grammatically marked) to begin with.
@@
Section 3 will look at both examples in more detail, Section 5 will apply the model to a classic XAI example.
@@ 2 Aspects of Understanding: Forms of knowledge
For our model of understanding, we follow primarily two conceptualizations: the types and qualities of knowledge as described in detail by de Jong and Ferguson-Hessler (1996), and the taxonomies of learning from the field of education (‘Bloom’s taxonomy’; Anderson and Krathwohl, 2001; Krathwohl et al., 1969).
de Jong and Ferguson-Hessler (1996) were interested in knowledge in instruction and learning and especially as a prerequisite for problem solving. They distinguished and elaborated characteristics of different types of knowledge. In terms of understanding as a goal and the achievement of explanations, two types of knowledge are important: conceptual knowledge is “knowledge of facts, concepts, and principles” (p. 107), whereas procedural knowledge refers to “actions and manipulations” that transform one state into another. Thus, in our terminology, conceptual knowledge describes ‘knowing-that’, and procedural knowledge refers to ‘knowing-how’. In the context of XAI, we focus on cognitive forms of procedural knowledge, such as written division and programming. Actions and skills with a higher motor component, such as cycling or knitting, may require other forms of more embodied explanations (showing).
@@
Building on the taxonomy of qualities of knowledge also developed by de Jong and Ferguson-Hessler (1996), we consider the following characteristics to be relevant for describing understanding:
@@
Both conceptual and procedural knowledge are described as more or less shallow or deep. Regarding conceptual knowledge, de Jong and Ferguson-Hessler (1996, p. 111) illustrate the distinction with symbols or formulas on a shallow level vs. concepts and their relations on a deep level. Thus, deep comprehension most likely reflects highly interconnected and elaborated concepts. Shallow comprehension would correspond to little interconnectedness, but also to few concepts. In terms of procedural knowledge, there are rules or recipes at a shallow level in contrast to meaningful actions at a deep level.
Explainees may require different levels of understanding, and explanations should be tailored to those needs (Wilson and Keil, 2000). Thus, it is not always necessary – or even annoying – to give or receive a long and complex explanation that leads to deep understanding. Often the explainer is satisfied with a single action or a simple fact.
@@
Automation of knowledge is particularly important for understanding and therefore for explanation. Consider how difficult it is to explain simple skills like riding a bicycle. This is because there is a strong distinction between declarative conceptual knowledge and non-declarative compiled procedural knowledge. Compilation enables automatic processes. The term “knowledge compilation”, inspired by computer science, is used by Anderson (1983) to describe how multiple cognitive steps are combined into a single routine. Compilation involves two stages: composition, where steps are combined, and proceduralization, where declarative knowledge becomes automatic, resulting in more fluid performance.
We conceive comprehension as involving conceptual knowledge, i.e., remembering facts about and experiences with various concepts such as theories and/or related routine actions. In contrast, enabledness is conceived as the ability to perform routine and/or complex actions (mentally and physically) in a competent and flexible manner. Furthermore, whereas comprehension is always ‘explicit knowledge’ that can be acquired through books or explicit verbal instruction, enabledness consists at least in part of ‘tacit knowledge’ that is not easily shared (Lim and Ahmed, 2000).
However, conceptual knowledge can also be compiled, e.g., rote learning of definitions or mnemonic sentences instead of deep comprehension. On the other hand, deep enabledness requires declarative, mostly deep comprehension. In addition to implicit learning of procedures as a basis for skills, e.g., native language grammar, there is explicit learning, e.g., adult learning of a foreign language. Explaining how to drive a car is a good example of how procedures are first explained in a declarative way, and then the steps are proceduralized and composed over the years (Anderson, 1983). de Jong and Ferguson-Hessler (1996) explain automation by mentioning tacit knowledge that cannot be expressed. In order to give an explanation, the explainer needs non-tacit knowledge, also of procedures. In the following sections, this intertwining of understanding quality and type will be discussed in detail.
@@
Regarding modality, we added embodiment to the verbal and pictorial aspects given by de Jong and Ferguson-Hessler (1996). Embodied knowledge is conceptually similar to procedural knowledge (Tanaka, 2011, p. 150), and it can be better performed than explained verbally. Embodiment plays an important role in learning situations where new members of a (scientific) community are taught to become skilled practitioners. Visibly embodied displays (e.g., how to hold a pencil, how to hold a Munsell color chart next to a piece of dirt to describe its color) are part of the knowledge conveyed in such situations (Goodwin, 2013).
@@
In addition to this characterization of different types of knowledge and their manifold qualities, as a further conceptualization, taxonomies of learning objectives help to describe the types and levels of understanding (e.g., Krathwohl et al., 1969; Anderson and Krathwohl, 2001). Anderson and Krathwohl 2001 (summarized in Anderson 2002) systematize educational objectives in two dimensions, the structure of the knowledge and the cognitive processes that are to be done with the knowledge.
@@
The structure of knowledge aims at a differentiation of the above mentioned ‘types of knowledge’. It includes factual, conceptual, procedural, and metacognitive knowledge. While factual knowledge contains the basic elements of a domain (e.g., definitions), conceptual knowledge contains the structure of the basic elements, e.g., principles, models, and theories. Comparable to de Jong and Ferguson-Hessler (1996), procedural knowledge aims at ‘how to do something’, such as skills and algorithms. The fourth form of knowledge, metacognition, will be neglected in our paper, because it is only of interest in special explanations regarding (own) cognition. Thus, the categories of knowledge are similar to the types of knowledge discussed by de Jong and Ferguson-Hessler (1996) with the additional hint that there are very basic elements of knowledge. In terms of understanding, factual and conceptual knowledge aim at comprehension, while procedural knowledge, of course, aims at enabledness.
@@
The second dimension of the taxonomy of learning aims at increasing the levels of cognitive processes that can be done with knowledge, namely remembering, understanding (note that this is a different use of the term), applying, analyzing, evaluating, and creating (Chi and Wylie, 2014). This strengthens the aspect of the level of knowledge introduced by de Jong and Ferguson-Hessler (1996). There is a gain in complexity from low level cognitive processes like remembering to applying to high level processes like creating something new – resulting in a hierarchy from shallow to deep understanding. However, assessing the level of understanding makes it necessary to take a closer look at these processes. For example, whether creating something new is a high-level process could depend on what is meant by ‘new’ and the complexity of what is being created, e.g., a new sentence or a scientific article.
@@
Figure 1: Forms of understanding as a four-field diagram. Understanding can be of the type comprehension and/or enabledness, each of which has two levels: shallow and deep. The four fields are based on the possible combinations: e.g., shallow comprehension and shallow enabledness (top left, Section 3.1), shallow comprehension and deep enabledness (top right, Section 3.2), deep comprehension and shallow enabledness (bottom left, Section 3.3), and deep comprehension and deep enabledness (bottom left, Section 3.4).
@@
In summary, there are two forms of understanding – comprehension and enabledness. Both forms of understanding can be on different levels, from very shallow to very deep understanding. The four resulting combinations (cf. Figure 1) are intermediate steps or final achievements of explanations with the possible goal of agency, the ability to control and adapt the (digital) environment (Passey et al. 2018; Section 3.4). The combinations will be characterized in the next section. On this basis, (in Section 4) we explain dynamics of comprehension in explanations, taking into account two relevant aspects: We consider differences between novices and experts in explanations and discuss how in the course of an explanation comprehension and enabledness are intertwined.